OpenStudy (anonymous):

Two plane mirrors are inclined at an angle (theta).if a light ray comes parallel to one mirror and strikes another thus finally emerges parallel to the mirror it striked first. Find theta?

OpenStudy (anonymous):

45degres

OpenStudy (anonymous):

No its not

OpenStudy (anonymous):

 you can see that \[180 - 2\theta = \theta\] so \[\theta = 60\deg\]